OpenStudy (lgbasallote):

the driver of a car slams on the brakes when he sees a tree blocking the road. the car slows significantly with an acceleration of -5.60 m/s^2 for 4.20 s, making straight skid marks 62.4 m long all the way to the tree. with what speed does the car then strike the tree? @mathslover i got 10 something but my book says 3 something...how? o.O

mathslover (mathslover):

the book is correct @lgbasallote

OpenStudy (kropot72):

I got an answer of 3.1 m/s thus: Assuming constant deceleration \[\frac{v _{2}-v _{1}}{t _{2}-t _{1}}=\frac{v-v _{0}}{t-0}\] \[-5.6=\frac{v-v _{0}}{4.2}...............(1)\] Average speed = 62.4/4.2 = speed after 2.1 seconds (4.2/2 seconds) Speed at t(0) = 14.857 + at = 14.857 + (5.6 * 2.1) = 26.617 m/s \[v _{0}=26.617 m/s\] Substituting the value for v(0) into equation (1) gives \[-(5.6\times 4.2)=v-26.617\] v = 3.1 m/s

mathslover (mathslover):

here is given , a = -5.60 m/s^2 , t = 4.2 seconds and s = 62.4 m since s = ut +\(\frac{1}{2}\)at^2 hence we will put the values of s , a and t here : \[62.4 = u(4.2)+\frac{1}{2}(-5.60)(4.2)^2\] \[62.4=4.2 u +49.392\] \[62.4-49.392-4.2u\] \[13.008=4.2u\] \[u = \frac{13.008}{4/2}\] \[u = 3 (approx.)\]

OpenStudy (aravindg):

ami late?

mathslover (mathslover):

@lgbasallote have u got that solution

OpenStudy (lgbasallote):

why + 49.392? shouldnt it be negative??

OpenStudy (lgbasallote):

@mathslover

mathslover (mathslover):

wait for 1 minute @lgbasallote i think i did mistake

mathslover (mathslover):

wait this is the correct solution now \[s=ut+\frac{1}{2}at^2\] \[62.4 = 4.2u -49.392\] \[62.4+49.392=4.2u\] \[111.792=4.2u\] \[26.6=u\] hence initial velocity = 26.6 m/s now final velocity : u + at therefore \(\large{v_f=v_i+at}\) that means \(\large{v_f=26.6 +(-5.6*4.2)}\) hence \(\huge{v_f=26.6-23.52}\) \(\huge{v_f=3.08 = 3.1ms^{-1}}\)

mathslover (mathslover):

@lgbasallote did u get it now ?

OpenStudy (lgbasallote):

i mistakenly wrote 6.24 in my solution :O no wonder...

mathslover (mathslover):

:) np . well sorry for that mistake